# Decode Neural signal as Speech

### v2

@@ Abstract
Decoding language from brain dynamics is an important open direction in the realm of brain-computer interface (BCI), especially considering the rapid growth of large language models. Compared to invasive-based signals which require electrode implantation surgery, non-invasive neural signals (e.g. EEG, MEG) have attracted increasing attention considering their safety and generality. However, the exploration is not adequate in three aspects: 1) previous methods mainly focus on EEG but none of the previous works address this problem on MEG with better signal quality; 2) prior works have predominantly used “teacher-forcing” during generative decoding, which is impractical; 3) prior works are mostly “BART-based” not fully auto-regressive, which performs better in other sequence tasks. In this paper, we explore the brain-to-text translation of MEG signals in a speech-decoding formation. Here we are the first to investigate a cross-attention-based “whisper” model for generating text directly from MEG signals without teacher forcing. Our model achieves impressive BLEU-1 scores of 60.30 and 52.89 without pretraining & teacher-forcing on two major datasets (GWilliams and Schoffelen). This paper conducts a comprehensive review to understand how speech decoding formation performs on the neural decoding tasks, including pretraining initialization, training & evaluation set splitting, augmentation, and scaling law. Code is available at https://github.com/NeuSpeech/NeuSpeech1/tree/main.
@@ 1 Introduction
Decoding language from brain activity is an important and rapidly advancing area of neurotechnology. The ability to infer a person’s intended words, phrases, or even full sentences directly from their brain signals has profound applications for improving communication and control capabilities for those suffering from severe speech and motor impairments. Conditions like high-level spinal cord injury or late-stage ALS can trap a person inside their own body, unable to fully express themselves or interact with the digital world around them. By noninvasively measuring and modeling the complex neural representations underlying language production in the brain, we aim to restore digital communication simply through thought. Beyond assisting those with disabilities, proficient decoding of full linguistic expressions from brain signals could also enable entirely new modes of seamless human-machine interaction. One could silently yet precisely command prosthetic devices, control software applications, browse the internet, or navigate virtual and augmented reality environments entirely using their brain. The potential applications seem nearly limitless and could vastly expand what is possible for both able-bodied individuals and those with severe impairments. It is an enormously challenging task from an artificial intelligence and neuroscience perspective but one with immense promise to transform lives if achieved.
@@
Notably, using decoding speech on invasive signals achieved high-performance recognition accuracy for real-time translation (Metzger et al., 2023; Willett et al., 2023). However, the invasive recordings pose medical risks and issues with maintaining over long periods. Thus, for non-invasive brain recordings, Wang et al. (Wang & Ji, 2022) demonstrated open-vocabulary EEG-to-text translation by employing pre-trained language models on word-level EEG features. Duan et al. (Duan et al., 2023) improved this approach by decoding from raw EEG waves without using time markers. However, limitations remain as evaluation relies on teacher forcing and the model cannot generate meaningful sentences without it. On MEG, previous works can merely decode limited classes from MEG signal. (Dash et al., 2020; Csaky et al., 2023; Ghazaryan et al., 2023). Recently, Defossez et al.(Défossez et al., 2023) decoded speech perception of 3 second segments from MEG signals, and was capable of classifying sentences by matching MEG signal with speech. However, it cannot generate sentences from the MEG signal.
@@
To address the issue of teacher forcing evaluation and generating text from MEG signal. In this paper, we present NeuSpeech, a pioneering framework that decodes neural as speech. NeuSpeech uses an encoder-decoder model (whisper (Radford et al., 2023)) rather than BART-based models that are widely used in previous works (Wang & Ji, 2022; Duan et al., 2023; Xi et al., 2023) to directly translate raw MEG waves into text, without pretraining or transforming neural signals to discrete codes. This approach 1) lets the model efficiently learn text-related information from neural signals, and 2) keeps the general meaning of the whole sentence when generating long texts.
@@
Experiments employ non-invasive MEG signals from the large-scale public GWilliams (Gwilliams et al., 2023) and Schoffelen (Schoffelen et al., 2019) datasets, which recorded MEG during speech listening tasks. Notably, NeuSpeech can generalize to all languages and equipment types. Performance is assessed using standard translation metrics (Papineni et al., 2002; Lin, 2004). For raw MEG waves without event markers, NeuSpeech achieves 60.30 BLEU-1 and 55.26 Rouge-1 F-measure on GWilliams, and 53.16 BLEU-1 on Schoffelen without teacher forcing evaluation. The main contributions of this work are as follows:
  • NeuSpeech adoptes an end-to-end neural framework for open-vocabulary MEG-to-text translation without markers without pre-training or teacher forcing, and is the first to achieve high-performance raw MEG wave-to-text translation.
  • We explore multi-layout, multi-language, multi-subject MEG-to-text transcription on multiple datasets under various training settings. This provides a more comprehensive evaluation than prior work.
  • We exploit cross-attention to get global semantic information for sentences rather than previous token-based methods.
2 Related Works
Significant progress has been made in the field of brain-to-text decoding in recent years. In 2019, Anumanchipalli et al. (Anumanchipalli et al., 2019) proposed a model that could decode ECoG signals into the kinematics of articulation, followed by acoustic features (such as MFCCs), and ultimately generate the speech waveform, which inspired many researchers to explore this area further. In 2020, Wang et al. (Wang et al., 2020) introduced the use of generative adversarial networks (GANs) to decode ECoG signals into speech. In 2021, Willett et al. (Willett et al., 2021) proposed a model that utilized an RNN and a probabilistic language model to decode letters while a subject was performing handwriting. Metzger et al. (Metzger et al., 2022) developed a pipeline that decoded ECoG signals into text using an RNN and then refined the results using the GPT-2 model. In the domain of open-vocabulary decoding, Metzger et al. (Metzger et al., 2023) proposed an RNN framework that could decode speech, text, emotion, and facial expressions from ECoG signals in real-time. Simultaneously, Willett et al. (Willett et al., 2023) successfully decoded text from brain signals. Liu et al. (Liu et al., 2023) developed a three-part model to decode logosyllabic languages (e.g., Chinese) by first decoding ECoG signals into Chinese pinyin with tones and syllables, and then synthesizing speech. Feng et al. (Feng et al., 2023) also achieved text decoding from SEEG signals. However, it is important to note that these practical systems predominantly rely on invasive brain recordings.
For non-invasive recordings, Meta proposed a brain-to-speech framework using contrastive learning with MEG and EEG signals (Défossez et al., 2023), which is capable of classifying limited sentences but cannot be used for open-vocabulary text decoding. Ghazaryan et al. (Ghazaryan et al., 2023) attempted to decode limited words using MEG responses. Wang et al. (Wang & Ji, 2022) developed a framework to decode word-level EEG signal features into text using a pre-trained BART model (Lewis et al., 2020). Many subsequent works have followed in the footsteps of Wang et al. (Wang & Ji, 2022), including Dewave (Duan et al., 2023), which proposed a framework that employed wave2vec (Baevski et al., 2020), and discrete codex to obtain effective representations and fed them into a BART (Lewis et al., 2020) model to generate text. However, these methods rely on teacher forcing, which is not feasible in real-world settings. Different from these methods, our work uses an end-to-end paradigm to generate open-vocabulary text and evaluate it without teacher forcing.
Table 1: Evaluation metrics of MEG-to-Text translation under raw waves input. For a fair comparison, these results has w/ and w/o the teacher-forcing evaluation setting since EEG-to-Text (Wang & Ji, 2022) only has teacher-forcing originally. Default is w/o teacher forcing and ‘w/ tf’ means with teacher-forcing.
BLEU-N (%) $\uparrow$ ROUGE-1 (%$\uparrow$ WER (%) $\uparrow$
Modality Method N=1 N=2 N=3 N=4 F P R
Speech Whisper (Radford et al., 2023) 67.79 59.15 51.97 45.78 75.90 73.53 81.24 41.21
Noise NeuSpeech 0.0 0.0 0.0 0.0 2.84 23.9 1.5 99.98
- random 4.99 1.18 0.59 0.41 4.79 5.95 5.87 132.30
MEG eeg-to-text (Wang & Ji, 2022) 9.21 2.13 0.57 0.14 9.74 10.73 11.38 118.25
MEG NeuSpeech 57.68 52.76 48.84 45.42 60.24 61.86 61.92 58.89
MEG NeuSpeech w/ pp 60.3 55.26 51.24 47.78 58.73 60.88 59.76 56.63
MEG eeg-to-text (Wang & Ji, 2022) w/ tf 21.66 9.27 3.93 1.73 20.11 22.48 18.71 95.69
MEG NeuSpeech w/ tf 78.13 71.25 64.79 58.86 81.91 80.52 83.68 25.38
MEG NeuSpeech w/ pp tf 80.20 73.27 66.79 60.81 78.62 78.14 79.54 23.42
@@ 3 Method
NeuSpeech is a modification of whisper (Radford et al., 2023) model, where the raw MEG signals are sent into the model and generate text. Whisper (Radford et al., 2023) is the most powerful and noise-robust model in speech recognition which is trained on hundreds of thousands of audio data and can transcribe speech from most of the languages on earth.
@@ 3.1 Task Definition
Given a raw MEG waves $\mathcal{X}$, the aim is to decode the corresponding open-vocabulary text tokens $\mathcal{W}$. These MEG-Text pairs $\langle\mathcal{X},\mathcal{W}\rangle$ are collected during hearing, as defined in Section 4.1. We decode neural signals as speech, and we only have one setting which is: Raw MEG Waves to Text Translation, where MEG waves $\mathcal{X}$ are directly sent into the whisper model to generate continuous words without teacher forcing, a more challenging but practical setting. NeuSpeech is the pioneering work in this task. Different from previous works(Duan et al., 2023; Wang & Ji, 2022), we are the first to evaluate generated text without teacher forcing using MEG.
@@ 3.2 Preprocessing
We only implement basic pre-processing here, the MEG signals are notched at power line frequency and filtered with (1,60) Hz. Then, both of them are resampled to 200Hz. After that, the signals are normalized using robust scaler mentioned in (Défossez et al., 2023), then clipped to ±$\pm 10$ and scaled into a value range of $(-1,1)$.
@@ 3.3 Model
Different from previous works (Duan et al., 2023; Tang et al., 2022) which are usually trained through multi-steps. Our model is trained end-to-end from scratch without bells and whistles. We freeze parameters in the decoder and only train the encoder. Since the original input channel of the whisper model is incompatible with the MEG channels, we replace the first convolution layer of the model with two convolution layers to fit in MEG channels and align time width respectively. For training data, we do not use any data augmentation. In terms of efficiency, we use Adalora (Zhang et al., 2023) technique to tune our model, and only the encoder including convolution layers are in the Lora model.
@@ 3.4 Evaluation
We evaluate transcribing performance on GWilliams dataset (Gwilliams et al., 2023) and Schoffelen (Schoffelen et al., 2019) using NLP metrics, BLEU, ROUGE and WER, which are used in previous works (Wang & Ji, 2022; Duan et al., 2023). We add three baselines to give a rough understanding of up and down limits. 1) speech recognition, we use a pre-trained whisper base model to demonstrate the high performance on speech modality. 2) input noise, we input Gaussian noise into our model to check if our model can distinguish between notoriously noisy MEG and noise, and the well-trained model should not generate text with pure noise as input. 3) random selection, we also showed the performance of randomly selecting sentences from training sets as predictions to give the down limit of the dataset as learned models should pass this limit to prove its effectiveness.
@@ 4.1 Dataset
We utilize GWilliams (Gwilliams et al., 2023) and Schoffelen (Schoffelen et al., 2019) datasets which are auditory MEG task for experiments. Both have a detailed timetable for each sentence, and they are divided by sentence-recording pairs into train:validation:test 8:1:1, which results in 23339:2917:2918 and 8570:1112:1076 MEG-text pairs respectively. However, we should note that the unique sentences are 661 and 360 for two datasets, which are very limited compared to speech datasets.
Table 2: Transcription results on the unseen MEG waves, where bold denotes a correct match between ground truth and our prediction. Underline denotes a fuzzy match with similar semantic meanings. For a fair comparison, these results do not keep the teacher-forcing evaluation setting as EEG-to-Text (Wang & Ji, 2022) and Dewave (Duan et al., 2023). This means the decoding process can hold on accumulated errors and predicts the next token with model-generated tokens from previous steps.
Decoding Results on GWilliams (Gwilliams et al., 2023)

### v3

# NeuSpeech: Decode Neural signal as Speech

@@ Abstract
Decoding language from brain dynamics is an important open direction in the realm of brain-computer interface (BCI), especially considering the rapid growth of large language models. Compared to invasive-based signals which require electrode implantation surgery, non-invasive neural signals (e.g. EEG, MEG) have attracted increasing attention considering their safety and generality. However, the exploration is not adequate in three aspects: 1) previous methods mainly focus on EEG but none of the previous works address this problem on MEG with better signal quality; 2) prior works have predominantly used “teacher-forcing" during generative decoding, which is impractical; 3) prior works are mostly “BART-based" not fully auto-regressive, which performs better in other sequence tasks. In this paper, we explore the brain-to-text translation of MEG signals in a speech-decoding formation. Here we are the first to investigate a cross-attention-based “whisper" model for generating text directly from MEG signals without teacher forcing. Our model achieves impressive BLEU-1 scores of 60.30 and 52.89 without pretraining & teacher-forcing on two major datasets (GWilliams and Schoffelen). This paper conducts a comprehensive review to understand how speech decoding formation performs on the neural decoding tasks, including pretraining initialization, training & evaluation set splitting, augmentation, and scaling law. Code is available at https://github.com/NeuSpeech/NeuSpeech1.
11footnotetext: The Hong Kong University of Science and Technology (Guangzhou), People’s Republic of China,
Yiqian Yang and Hui Xiong are with AI Thrust, HKUST(GZ).
Qiang Zhang and Renjing Xu are with MICS Thrust, HKUST(GZ),
Jinni Zhou is with RBM Base, College of Future Technology,
Email: yyang937@connect.hkust-gz.edu.cn, xionghui@hkust-gz.edu.cn, qzhang749@connect.hkust-gz.edu.cn, renjingxu@hkust-gz.edu.cn, eejinni@hkust-gz.edu.cn
22footnotetext: GrapheneX-UTS HAI Centre, Australia Artificial Intelligence Institute, University of Technology Sydney, Australia,
Email:duanyiquncc@gmail.com
33footnotetext: Department of Software Convergence, Kyung Hee University, Republic of Korea,
Email:girlsending0@khu.ac.kr, whlee@khu.ac.kr
@@ 1 Introduction
Decoding language from brain activity is an area of neurotechnology that’s rapidly evolving, offering significant benefits for improving communication and control for people with severe speech and motor impairments. Individuals with conditions like high-level spinal cord injuries or late-stage ALS often face significant communication challenges. Through noninvasive measurement and modeling of neural representations related to language production, we aim to enable digital communication based on thoughts alone. This advancement has the potential to transform the lives of people with disabilities, providing them with seamless and intuitive communication and device control capabilities.
@@
Notably, using decoding speech on invasive signals achieved high-performance recognition accuracy for real-time translation [1, 2]. However, the invasive recordings pose medical risks and issues with maintaining over long periods. Thus, for non-invasive brain recordings, Wang et al. [3] demonstrated open-vocabulary EEG-to-text translation by employing pre-trained language models on word-level EEG features. Duan et al. [4] improved this approach by decoding from raw EEG waves without using time markers. However, limitations remain as evaluation relies on teacher forcing and the model cannot generate meaningful sentences without it. On MEG, previous works can merely decode limited classes from MEG signal. [5, 6, 7]. Recently, Defossez et al.[8] decoded speech perception of 3 second segments from MEG signals, and was capable of classifying sentences by matching MEG signal with speech. However, it cannot generate sentences from the MEG signal.
@@
To address the issue of teacher forcing evaluation and generating text from MEG signal. In this paper, we present NeuSpeech, a pioneering framework that decodes neural as speech. NeuSpeech uses an encoder-decoder model (whisper [9]) rather than BART-based models that are widely used in previous works [3, 4, 10] to directly translate raw MEG waves into text, without pretraining or transforming neural signals to discrete codes. This approach 1) lets the model efficiently learn text-related information from neural signals, and 2) keeps the general meaning of the whole sentence when generating long texts.
Experiments employ non-invasive MEG signals from the large-scale public GWilliams [11] and Schoffelen [12] datasets, which recorded MEG during speech listening tasks. Notably, NeuSpeech can generalize to all languages and equipment types. Performance is assessed using standard translation metrics [13, 14]. For raw MEG waves without event markers, NeuSpeech achieves 60.30 BLEU-1 and 55.26 Rouge-1 F-measure on GWilliams, and 53.16 BLEU-1 on Schoffelen without teacher forcing evaluation. Decoding language from brain activity is an important and rapidly advancing area of neurotechnology. The ability to infer a person’s intended words, phrases, or even full sentences directly from their brain signals has profound applications for improving communication and control capabilities for those suffering from severe speech and motor impairments. Conditions like high-level spinal cord injury or late-stage ALS can trap a person inside their own body, unable to fully express themselves or interact with the digital world around them. By non-invasively measuring and modeling the complex neural representations underlying language production in the brain, we aim to restore digital communication simply through thought. Beyond assisting those with disabilities, proficient decoding of full linguistic expressions from brain signals could also enable entirely new modes of seamless human-machine interaction. One could silently yet precisely command prosthetic devices, control software applications, browse the internet, or navigate virtual and augmented reality environments entirely using their brain. The potential applications seem nearly limitless and could vastly expand what is possible for both able-bodied individuals and those with severe impairments. It is an enormously challenging task from an artificial intelligence and neuroscience perspective but one with immense promise to transform lives if achieved.
Notably, using decoding speech on invasive signals achieved high-performance recognition accuracy for real-time translation [1, 2]. However, the invasive recordings pose medical risks and issues with maintaining over long periods. Thus, for non-invasive brain recordings, Wang et al. [3] demonstrated open-vocabulary EEG-to-text translation by employing pre-trained language models on word-level EEG features. Duan et al. [4] improved this approach by decoding from raw EEG waves without using time markers. However, limitations remain as evaluation relies on teacher forcing and the model cannot generate meaningful sentences without it. On MEG, previous works can merely decode limited classes from MEG signal. [5, 6, 7]. Recently, Defossez et al.[8] decoded speech perception of 3 second segments from MEG signals, and was capable of classifying sentences by matching MEG signal with speech. However, it cannot generate sentences from the MEG signal.
To address the issue of teacher forcing evaluation and generating text from MEG signal. In this paper, we present NeuSpeech, a pioneering framework that decodes neural as speech. NeuSpeech uses an encoder-decoder model (whisper [9]) rather than BART-based models that are widely used in previous works [3, 4, 10] to directly translate raw MEG waves into text, without pretraining or transforming neural signals to discrete codes. This approach 1) lets the model efficiently learn text-related information from neural signals, and 2) keeps the general meaning of the whole sentence when generating long texts.
@@
Experiments employ non-invasive MEG signals from the large-scale public GWilliams [11] and Schoffelen [12] datasets, which recorded MEG during speech listening tasks. Notably, NeuSpeech can generalize to all languages and equipment types. Performance is assessed using standard translation metrics [13, 14]. For raw MEG waves without event markers, NeuSpeech achieves 60.30 BLEU-1 and 55.26 Rouge-1 F-measure on GWilliams, and 53.16 BLEU-1 on Schoffelen without teacher forcing evaluation.
@@ 2 Method
NeuSpeech is a modification of whisper [9] model, where the raw MEG signals are sent into the model and generate text. Whisper [9] is the most powerful and noise-robust model in speech recognition which is trained on hundreds of thousands of audio data and can transcribe speech from most of the languages on earth.
@@ 2.1 Task Definition
Given a raw MEG waves $\mathcal{X}$, the aim is to decode the corresponding open-vocabulary text tokens $\mathcal{W}$. These MEG-Text pairs $\langle\mathcal{X},\mathcal{W}\rangle$ are collected during hearing, as defined in Sec. 3.1. We decode neural signals as speech, and we only have one setting which is: Raw MEG Waves to Text Translation, where MEG waves $\mathcal{X}$ are directly sent into the whisper model to generate continuous words without teacher forcing, a more challenging but practical setting. NeuSpeech is the pioneering work in this task. Different from previous works[4, 3], we are the first to evaluate generated text without teacher forcing using MEG.
@@ 2.2 Preprocessing
We only implement basic pre-processing here, the MEG signals are notched at power line frequency and band-pass filtered within 1 Hz and 60 Hz. Then, it is resampled to 200Hz. After that, the signal is normalized using robust scaler mentioned in [8], then clipped to ±$\pm 10$ and scaled into a value range of $(-1,1)$.
@@ 2.3 Model
Figure 1: NeuSpeech overview. MEG signal is recorded while the subject is listening to speech. Our model is trained in an end-to-end manner, and only trains the AdaLora module applied on encoder and convolution layers using cross-entropy loss. In evaluation, we tested in both situations, w/ and w/o teacher forcing. In testing, teacher forcing means predict next token using previous ground truth tokens, rather than model generated tokens.
Different from previous works [4, 15] which are usually trained through multi-steps. Our model is trained end-to-end from scratch without bells and whistles. We freeze parameters in the decoder and only train the encoder. Since the original input channel of the whisper model is incompatible with the MEG channels, we replace the first convolution layer of the model with two convolution layers to fit in MEG channels and align time width respectively. For training data, we do not use any data augmentation. In terms of efficiency, we use Adalora [16] technique to tune our model, and only the encoder including convolution layers are in the Lora model.
@@ 2.4 Evaluation
We evaluate transcribing performance on GWilliams dataset [11] and Schoffelen [12] using NLP metrics, BLEU, ROUGE and WER, which are used in previous works [3, 4]. We add three baselines to give a rough understanding of up and down limits. 1) speech recognition, we use a pre-trained whisper base model to demonstrate the high performance on speech modality. 2) input noise, we input Gaussian noise into our model to check if our model can distinguish between notoriously noisy MEG and noise, and the well-trained model should not generate text with pure noise as input.
@@ 3.1 Dataset
We utilize GWilliams [11] and Schoffelen [12] datasets which are auditory MEG task for experiments. We only used auditory sentence data in Schoffelen. Both datasets have a detailed timetable for each sentence, and they are divided by sentence-recording pairs into train:validation:test 8:1:1, which results in 23339:2917:2918 and 8570:1112:1076 MEG-text pairs respectively. Unique words in GWilliams and Schoffelen is 2260 and 1861 respectively. We should note that the unique sentences are 661 and 360 for two datasets, which are very limited compared to speech datasets. More details at Supp. A.
3.2 Implementation Details
Model modification. We adopt whisper base model [9] as our base model. To adapt the input size of the neural signal, we use two convolution1d layers, the kernel sizes are 3, and stride is 1,2 respectively to replace the first convolution layer in the whisper-base model of dimension 512. Detailed Python code at supp.D.
Training. To implement eeg-to-text [3], we use the original word timestamps provided by GWilliams [11] to segment MEG waves for each word feature. The training settings are the same as the origin code.
All models are trained on Nvidia H800(80GB) GPUs. We use a learning rate of 1e-3 and batch size of 64 for training 500 epochs and select the best model on evaluation loss. We use the AdamW as the optimizer for training all the models. We used early stopping on evaluation loss, if the evaluation loss does not decrease in 5 epochs, the training is stopped. It takes about 48 hours to train one model on single GPU.
Evaluation. For non-teacher forcing generation configuration, we use the beam search with a beam size of 5 and set the repetition penalty to 5, with no repeat n-gram size to 2, which is very important to avoid repeating.
3.3 Metrics
Table 1: Evaluation metrics of MEG-to-Text translation under raw waves input. For a fair comparison, these results has w/ and w/o the teacher-forcing evaluation setting since EEG-to-Text [3] only has teacher-forcing originally. Default is w/o teacher forcing and ‘w/ tf’ means with teacher-forcing.
BLEU-N (%) $\uparrow$ ROUGE-1 (%$\uparrow$ WER (%) $\downarrow$
Modality Method N=1 N=2 N=3 N=4 F P R
Speech Whisper [9] 67.79 59.15 51.97 45.78 75.90 73.53 81.24 41.21
Noise NeuSpeech 0.0 0.0 0.0 0.0 2.84 23.9 1.5 99.98
MEG eeg-to-text [3] 9.21 2.13 0.57 0.14 9.74 10.73 11.38 118.25
MEG NeuSpeech 57.68 52.76 48.84 45.42 60.24 61.86 61.92 58.89
MEG eeg-to-text [3] w/ tf 21.66 9.27 3.93 1.73 20.11 22.48 18.71 95.69
MEG NeuSpeech w/ tf 78.13 71.25 64.79 58.86 81.91 80.52 83.68 25.38
As shown in Table 1, we compare our results to input Gaussian noise to the NeuSpeech model and randomly select sentences from the test set and whisper base model speech transcription to give the understanding of up and down limits. We also compare previous method eeg-to-text [3].
Based on the information provided in Tab. 1, it is evident that our method outperforms randomly selected sentences, noise input, and eeg-to-text [3]. NeuSpeech has demonstrated the feasibility of decoding text from non-invasive brain recordings. Prior to NeuSpeech, many previous works [4, 3] in this field relied on teacher forcing to evaluate their models, Defossez et al. [8] used classification accuracy. However, NeuSpeech introduces a more fair and real-world setting by employing end-to-end training, which directly ties neural signals to text. In comparison to previous works evaluated using teacher forcing, NeuSpeech has shown remarkable improvement, achieving BLEU-1, BLEU-2, BLEU-3, and BLEU-4 scores of 60.3, 55.26, 51.24, and 47.78, respectively, which is close to the performance of speech recognition.
To ensure a fair comparison, we also implement teacher forcing for both the eeg-to-text model [3] and NeuSpeech. It is evident that our BLEU-1 score of 80.20 is almost four times that of the eeg-to-text [3] model, highlighting the superiority of our approach. Furthermore, we have surpassed speech recognition methods with teacher forcing, emphasizing that our model can significantly improve performance by eliminating accumulated errors. It is worth noting that, at this stage, BLEU is the most suitable metric for evaluating the MEG-to-text system, as ROUGE and WER tend to yield high scores for noise input models.
3.4 Generated Samples
Table 2: Transcription results on the unseen MEG waves, where bold denotes a correct match between ground truth and our prediction. Underline denotes a fuzzy match with similar semantic meanings. For a fair comparison, these results do not keep the teacher-forcing evaluation setting as EEG-to-Text [3] and Dewave [4]. This means the decoding process can hold on accumulated errors and predicts the next token with model-generated tokens from previous steps. More samples are provided in supp. E.
Decoding Results on GWilliams [11]